# Interval and $\ell$-interval rational parking functions

### v4

@@ 3.2. $\ell-$interval parking functions
$\left|\mathrm{\mathrm{IPF}}_{n}(\ell)\right|=\sum^{n-1}_{x=0}\binom{n-1}{x}\cdot\min\left(x+1,\ell+1\right)\cdot\left|\mathrm{\mathrm{IPF}}_{x}(\ell)\right|\cdot\left|\mathrm{\mathrm{IPF}}_{n-1-x}(\ell)\right|,$
@@
Table 1. Cardinalities of the set $\mathrm{\mathrm{IPF}}_{n}(\ell)$ for $0\leq n\leq 9$ and $0\leq\ell\leq 7$, added by the authors to OIES as sequence A365623.
3.3. $\ell-$interval rational parking functions
@@
Now we must sum over $10\leq k\leq z-$ 1 and over $1\leq z\leq m$ which yields the desired formula. ∎
4. Nondecreasing $\ell-$interval rational parking functions
@@
As before, we simplify notation and let $\mathrm{\mathrm{IPF}}_{n,n}^{\uparrow}(\ell)\coloneqq\mathrm{\mathrm{IPF}}_{n}^{\uparrow}(\ell)$ when we have the same number of cars as parking spots.
@@
Note that whenever $z_{i}=z_{i-1}+1$, then $|\mathrm{\mathrm{IPF}}_{z_{i}-{z_{i}-1}-1}^{\uparrow}(\ell)|=|\mathrm{\mathrm{IPF}}_{0}^{\uparrow}(\ell)|=1$ accounting for the empty preference list. ∎
@@
4.1. A connection to lattice paths
@@
Following [4], we recall that Dyck paths can be encoded via a Dyck word of length $2n$ consisting of $n$ letters $N$, representing north steps, and $n$ letters $E$, representing east steps, satisfying
  1. (1)
@@
Figure 4. Examples of two Dyck paths and their corresponding Dyck words.
@@
$j-a_{j}\geq\ell+1,$ (4.1)
as car $j$ with preference $a_{j}$ parked further than $\ell$ spots from its preference. By definition of $\varphi$, we know $a_{j}=1+|\{x\,:\,x<\mathrm{N}(j)\mbox{ and }w_{x}=E\}|$. In what follows, let |$|\{x\,:\,x<\mathrm{N}(j)\mbox{ and }w_{x}=E\}|=\#E$. Replacing $a_{j}$ with #$\#E+1$ in (4.1) yields
$j-\#E-1\geq\ell+1.$ (4.2)
At this point in $w$, we have taken $j$ many $N$ steps and the height of our Dyck path at step $j+\#E$ is #$\#N-\#E$. From inequality (4.2), this means that #$\#N-\#E-1\geq\ell+1$, which implies that #$\#N-\#E\geq\ell+2$. This contradicts that the height of this Dyck path is at most $\ell+1$. Therefore, $\varphi(w)\in\mathrm{\mathrm{IPF}}_{n}^{\uparrow}(\ell)$ as desired.
@@
For (iii). It now suffices to note that from part (1) in this proof, we have shown by contradiction that if $w$ does not correspond to a Dyck path with maximum height $\ell+1$, then $\varphi(w)$ cannot be in $\mathrm{\mathrm{IPF}}_{n}^{\uparrow}(\ell)$. ∎
The enumeration of Dyck paths based on their exact height is well-known (OEIS A080936) and is credited to Grossman [20], although we were unable to access the article to confirm this. Thanks to a comment from Emeric Deutsch posted on OEIS A080936 we found the following: By [25, Page 37-38] the number of Dyck paths with semilength $2n$ and height at most $k$, denoted $H(k)$, is given by
$H(k)=[t^{k}]\left(\frac{f_{k}(t)}{f_{k+1}(t)}\right),$ (4.4)
which is the coefficient of $t^{k}$ in the generating function $f_{k}(t)/f_{k+1}(t)$, where $f_{0}(t)=f_{1}(t)=1$ and $f_{k+1}(t)=f_{k}(t)-t\cdot f_{k-1}(t)$. Note that the recurrence relation for the polynomials $f_{k}(t)$ was given in [25, Page 36, Equation 14].
Given the bijection between nondecreasing $\ell$-interval parking functions of length $n$ and Dyck paths of semilength $2n$ with height at most $\ell+1$, and Equation (4.4), we arrive at the following result.
@@
For $k\geq 0$, the number of nondecreasing $\ell$-interval parking functions of length $n$ is $H(\ell+1)$, as defined in Equation (4.4).
We can extend this result naturally to $\ell$-interval rational parking functions, which are in bijection with rational $(n,m)$-Dyck paths (see [9] for an overview of rational Dyck paths and other Catalan objects, and see [11] for a relationship between $k$-Naples parking functions and $k$-lattice paths). Let us continue with some definitions and notation.
@@
Figure 5. A $(5,8)$-Dyck path with word NENNEEENENEEE with height 2.
Definition 4.12 (Pages 761-762 in [9]).
@@
To complete the proof it suffices to show that $\mathrm{\mathrm{IPF}}_{n}^{\uparrow}(\ell)\subseteq\varphi(\mathrm{D}_{n}^{\ell+1})$. The proof of this follows by an analogous argument as that presented in part (2) within the proof of Theorem 4.7. The only difference is that in the construction of the word $w$ in equation (4.3) there are $n$ north steps and $m$ east steps and the Dyck path has height at most $\ell+1$. ∎
@@
4.2. The members of $\mathrm{\mathrm{IPF}}_{n}^{\uparrow}(\ell)$ that are also Fubini rankings
We now consider the intersection of nondecreasing $\ell$-interval parking functions which are also Fubini rankings (Definition 2.3). The motivation for considering these sets is the recent work of Elder, Harris, Kretschmann, and Martínez Mori, who enumerated unit Fubini rankings which are unit interval parking functions of length $n$ that are also Fubini rankings. They established the following enumerative result [15, Theorem 3.3].
@@ 4.2. The members of $\mathrm{\mathrm{IPF}}_{n}^{\uparrow}(\ell)$ that are also Fubini rankings
$|\mathrm{FR}_{n}^{\uparrow}(\ell)|=\sum_{x=0}^{\ell}|\mathrm{FR}_{n-1-x}^{\uparrow}(\ell)|.\qed$
In Table 2, we give data for the cardinalities of the set $\mathrm{FR}_{n}^{\uparrow}(\ell)$ when $0\leq n\leq 11$ and $0\leq\ell\leq 6$. We direct interested readers to [26]).
@@
Table 2. Cardinalities of the set $\mathrm{FR}_{n}^{\uparrow}(\ell)$ for $0\leq n\leq 11$ and $0\leq\ell\leq 6$.
@@
As noted at the start of this section, Elder, Harris, Kretschmann, and Martínez Mori studied the set $\mathrm{FR}_{n}(1)$ [15]. They proved that $|\mathrm{FR}_{n}(1)|$ was equal to the number of Boolean intervals in the weak order lattice of $\mathfrak{S}_{n}$ [15, Theorem 1.2]. It remains an open problem to determine if there is some correspondence between the set $\mathrm{FR}_{n}(\ell)$ and other structures in the weak order lattice of $\mathfrak{S}_{n}$.
5. Unit interval rational parking functions and preferential arrangements
@@ 5.1. Preferential arrangements
Gross defines preferential arrangements in [19] as the number of ways $n$ distinguishable objects can be distributed into $r$ distinguishable boxes where $r$ varies $1\leq r\leq n$. Now recall that Fubini rankings are rankings of $n$ competitors in which ties are allowed (Definition 2.3). There is a bijection between preferential arrangements and Fubini rankings, which we prove next in Theorem 5.2.
@@
For surjectivity: Let $B_{1}B_{2}\cdots B_{r}\in\mathrm{PA}_{n,r}$. We construct a tuple in $[n]^{n}$ as follows: If $i\in B_{1}$, then set $a_{i}=1$. If $i\in B_{j}$ with $1<j\leq r$, then set $a_{i}=1+\sum_{k=1}^{j-1}|B_{k}|$. Let $\alpha=(a_{1},a_{2},\ldots,a_{n})$. Claim: $\alpha\in\mathrm{FR}_{n,r}$. First note that $\alpha$ has length $n$ as $\sum_{j=1}^{r}|B_{j}|=n$. By construction, the tuple $\alpha$ has precisely $r$ distinct values i.e., $r$ ranks. We need to show that the ranks in $\alpha$ satisfy the needed condition for a Fubini ranking. Note that there are $|B_{1}|$ entries in $\alpha$ with value one. Then the next rank appearing in $\alpha$ is precisely $1+|B_{1}|$, accounting for the fact that $|B_{1}|$ competitors tied for rank one. Note that then rank $1+|B_{2}|$ appears in $\alpha$ at index $i$ whenever $i\in B_{2}$. Continuing in this fashion, we then guarantee that all $r$ ranks are utilized and that no ranks are skipped. Thus $\alpha$ is a Fubini ranking, as claimed. It suffices to show that $\varphi(\alpha)=B_{1}B_{2}\cdots B_{r}$. By definition, all the elements of $\alpha$ in $I(1)$ will be placed in $B_{1}$. Similarly, if $x$ is the next available rank in $\alpha$, all the elements with rank $x$ will be placed in $B_{2}$, repeating this process for each of the $r$ available ranks ensures that $\varphi(\alpha)=B_{1}B_{2}\cdots B_{r}$, as claimed. ∎
@@
Proof.
Given that $\mathrm{PA}_{n}=\cup_{r=1}^{n}\mathrm{PA}_{n,r}$ and $\mathrm{FR}_{n}=\cup_{r=1}^{n}\mathrm{FR}_{n,r}$, and the fact that the function $\varphi$ in Theorem, 5.2 preserves the number of distinct ranks/boxes, we have that $\varphi:\mathrm{FR}_{n}\to\mathrm{PA}_{n}$ defined analogously as in Theorem 5.2 is also a bijection. ∎
@@
5.2. Unit interval rational parking functions and barred preferential arrangements
From [7], we know that the set of Fubini rankings are in bijection with the unit interval parking functions. In Theorem 5.2, we establish that Fubini rankings are also in bijection with preferential arrangements. Thus, preferential arrangements are in bijection with unit interval parking functions. In fact more is true. In this section, we establish that unit rational parking functions are in bijection with barred preferential arrangements.
Barred preferential arrangements were first introduced by Ahlbach et al. [3] and have been also studied by Nkonkobe and Murali [27] and they are a generalization of preferential arrangements, which we define next.
@@ 5.2. Unit interval rational parking functions and barred preferential arrangements
Definition 5.10 (Definition 2.7, [7]).
Given $\alpha=(a_{1},\ldots,a_{n})\in\mathrm{PF}_{n}$, let $\alpha^{\prime}=(a^{\prime}_{1},\ldots,a^{\prime}_{n})$ be the weakly increasing rearrangement of $\alpha$. The partition of $\alpha^{\prime}$ as a concatenation $\alpha^{\prime}=\pi_{1}|\pi_{2}|\cdots|\pi_{m}$, where $\pi_{j}$ begins at (and includes) the $j$th entry $\alpha^{\prime}_{i}$ satisfying $\alpha^{\prime}_{i}=i$, is called the block structure of $\alpha$. Each $\pi_{i}$ is called a block of $\alpha$.
@@
Corollary 5.11 (Corollary 3.1, [15]).
Let $\alpha\in\mathrm{UPF}_{n}$ and $b_{1}\,|\,b_{2}\,|\,\cdots\,|\,b_{m}$ be its block structure. For each $j\in[m]$, let $i_{j}$ be the minimal element of $b_{j}$. Consider any $j\in[m-1]$. If $|b_{j}|=1$, then $b_{j}=(i_{j})$ and $i_{j+1}=i_{j}+1$. Otherwise, if $|b_{j}|=2$, then $b_{j}=(i_{j},i_{j})$ and $i_{j+1}=i_{j}+2$. Otherwise, $|b_{j}|\geq 3$, $b_{j}=(i_{j},i_{j},\underbrace{i_{j}+1,i_{j}+2,\ldots,i_{j}+|b_{j}|-2}_{|b_{j}|-2\text{ terms }})$ and $i_{j+1}=i_{j}+|b_{j}|$.
@@
$a_{i}=\begin{cases}m_{j}&\mbox{if $i$ is the first entry in $B_{j}$}\\ m_{j}+y-2&\mbox{if $i$ is the $y$th entry in $B_{j}$}.\end{cases}$
@@
Now we put all the steps above together. Since $\mathcal{O}(\alpha)=x(q)$, steps 1 and 2 ensure we have a barred preferential arrangement in $\mathrm{BPA}(n,m-n)$. Then we add parenthesis satisfying $|\pi_{i}|=B_{i}$ for all $i\in[r]$ and so step 3 yields precisely the parenthesization of $q$. Therefore, $\vartheta(\alpha)=q$. ∎
We conclude with some enumerative consequences of Theorem 5.17 and the work of Ahlbach, Usatine, and Pippenger  [3].
@@
[3, Theorem 2] If $m\geq 1$, then
$|\mathrm{\mathrm{IPF}}_{n,m}(1)|=\frac{1}{2^{m-n}(m-n)!}\sum_{i=0}^{m-n}s(m-n+1,i+1)\thinspace|\mathrm{\mathrm{IPF}}_{n+i}(1)|$ (5.1)
@@
[3, Theorem 3] If $m\geq 0$ and $n\geq 1$, then
$|\mathrm{\mathrm{IPF}}_{n,m}(1)|=\sum_{k=0}^{n}k!\thinspace S(n,k)\thinspace\binom{m-n+k}{k}$ (5.2)
@@
Note that setting $m=n$ in Theorem 5.2 recovers Theorem 2.2. We conclude with the following illustrative example.
@@
Acknowledgements
This material is based upon work supported by the National Science Foundation under Grant No. DMS-1659138 and the Sloan Grant under Grant No. G-2020-12592. P. E. Harris was supported through a Karen Uhlenbeck EDGE Fellowship. J. Elder was partially supported through an AWM Mentoring Travel Grant. J. C. Martínez Mori is supported by Schmidt Science Fellows, in partnership with the Rhodes Trust. C. D. Monroe was partially supported by NSF GRFP Fellowship.

### v5

@@ 3.2 $\ell-$interval parking functions
$|\mathrm{\mathrm{IPF}}_{n}(\ell)|=\sum^{n-1}_{x=0}\binom{n-1}{x}\cdot\min\left(x+1,\ell+1\right)\cdot|\mathrm{\mathrm{IPF}}_{x}(\ell)|\cdot|\mathrm{\mathrm{IPF}}_{n-1-x}(\ell)|,$
@@
Table 1: Cardinalities of the set $\mathrm{\mathrm{IPF}}_{n}(\ell)$ for $0\leq n\leq 9$ and $0\leq\ell\leq 7$, added by the authors to OIES as sequence A365623.
3.3 $\ell-$interval rational parking functions
@@
Now we must sum over $10\leq k\leq z-$ 1 and over $1\leq z\leq m$ which yields the desired formula.
4 Nondecreasing $\ell-$interval rational parking functions
@@
As before, we simplify notation and let $\mathrm{\mathrm{IPF}}_{n,n}^{\uparrow}(\ell)=\mathrm{\mathrm{IPF}}_{n}^{\uparrow}(\ell)$ when we have the same number of cars as parking spots.
@@
Note that whenever $z_{i}=z_{i-1}+1$, then $|\mathrm{\mathrm{IPF}}_{z_{i}-{z_{i}-1}-1}^{\uparrow}(\ell)|=|\mathrm{\mathrm{IPF}}_{0}^{\uparrow}(\ell)|=1$ accounting for the empty preference list.
@@
4.1 A connection to lattice paths
@@
Following Armstrong et al. (2016), we recall that Dyck paths can be encoded via a Dyck word of length $2n$ consisting of $n$ letters $N$, representing north steps, and $n$ letters $E$, representing east steps, satisfying
  1. 1.
@@
Figure 4: Examples of two Dyck paths and their corresponding Dyck words.
@@
$j-a_{j}\geq\ell+1,$ (6)
as car $j$ with preference $a_{j}$ parked further than $\ell$ spots from its preference. By definition of $\varphi$, we know $a_{j}=1+|\{x\,:\,x<\mathrm{N}(j)\mbox{ and }w_{x}=E\}|$. In what follows, let |$|\{x\,:\,x<\mathrm{N}(j)\mbox{ and }w_{x}=E\}|=\#E$. Replacing $a_{j}$ with #$\#E+1$ in Equation (6) yields
$j-\#E-1\geq\ell+1.$ (7)
At this point in $w$, we have taken $j$ many $N$ steps and the height of our Dyck path at step $j+\#E$ is #$\#N-\#E$. From inequality (7), this means that #$\#N-\#E-1\geq\ell+1$, which implies that #$\#N-\#E\geq\ell+2$. This contradicts that the height of this Dyck path is at most $\ell+1$. Therefore, $\varphi(w)\in\mathrm{\mathrm{IPF}}_{n}^{\uparrow}(\ell)$ as desired.
@@
For (iii). It now suffices to note that from part (1) in this proof, we have shown by contradiction that if $w$ does not correspond to a Dyck path with maximum height $\ell+1$, then $\varphi(w)$ cannot be in $\mathrm{\mathrm{IPF}}_{n}^{\uparrow}(\ell)$.
The enumeration of Dyck paths based on their exact height is well-known (OEIS A080936) and is credited to Grossman (1950), although we were unable to access the article to confirm this. Thanks to a comment from Emeric Deutsch posted on OEIS A080936 we found the following: By (Kreweras, 1970b, Page 37-38) the number of Dyck paths with semilength $2n$ and height at most $k$, denoted $H(k)$, is given by
$H(k)=[t^{k}]\left(\frac{f_{k}(t)}{f_{k+1}(t)}\right),$ (9)
which is the coefficient of $t^{k}$ in the generating function $f_{k}(t)/f_{k+1}(t)$, where $f_{0}(t)=f_{1}(t)=1$ and $f_{k+1}(t)=f_{k}(t)-t\cdot f_{k-1}(t)$. Note that the recurrence relation for the polynomials $f_{k}(t)$ was given in (Kreweras, 1970b, Page 36, Equation 14).
Given the bijection between nondecreasing $\ell$-interval parking functions of length $n$ and Dyck paths of semilength $2n$ with height at most $\ell+1$, and Equation (9), we arrive at the following result.
@@
For $k\geq 0$, the number of nondecreasing $\ell$-interval parking functions of length $n$ is $H(\ell+1)$, as defined in Equation (9).
We can extend this result naturally to $\ell$-interval rational parking functions, which are in bijection with rational $(n,m)$-Dyck paths (see Ceballos and D’León (2019) for an overview of rational Dyck paths and other Catalan objects, and see Christensen et al. (2020) for a relationship between $k$-Naples parking functions and $k$-lattice paths). Let us continue with some definitions and notation.
@@
Figure 5: A $(5,8)$-Dyck path with word NENNEEENENEEE with height 2.
Definition 4.14 (Pages 761-762 in Ceballos and D’León (2019)).
@@
To complete the proof it suffices to show that $\mathrm{\mathrm{IPF}}_{n}^{\uparrow}(\ell)\subseteq\varphi(\mathrm{D}_{n}^{\ell+1})$. The proof of this follows by an analogous argument as that presented in part (2) within the proof of Theorem 4.8. The only difference is that in the construction of the word $w$ in Equation (8) there are $n$ north steps and $m$ east steps and the Dyck path has height at most $\ell+1$.
@@
4.2 The members of $\mathrm{\mathrm{IPF}}_{n}^{\uparrow}(\ell)$ that are also Fubini rankings
We now consider the intersection of nondecreasing $\ell$-interval parking functions which are also Fubini rankings (Definition 2.3). The motivation for considering these sets is the recent work of Elder, Harris, Kretschmann, and Martínez Mori, who enumerated unit Fubini rankings which are unit interval parking functions of length $n$ that are also Fubini rankings. They established the following enumerative result (Elder et al., 2023a, Theorem 3.3).
@@ 4.2 The members of $\mathrm{\mathrm{IPF}}_{n}^{\uparrow}(\ell)$ that are also Fubini rankings
$|\mathrm{FR}_{n}^{\uparrow}(\ell)|=\sum_{x=0}^{\ell}|\mathrm{FR}_{n-1-x}^{\uparrow}(\ell)|.$
In Table 2, we give data for the cardinalities of the set $\mathrm{FR}_{n}^{\uparrow}(\ell)$ when $0\leq n\leq 11$ and $0\leq\ell\leq 6$. We direct interested readers to Miles (1960)).
@@
Table 2: Cardinalities of the set $\mathrm{FR}_{n}^{\uparrow}(\ell)$ for $0\leq n\leq 11$ and $0\leq\ell\leq 6$.
@@
As noted at the start of this section, Elder, Harris, Kretschmann, and Martínez Mori studied the set $\mathrm{FR}_{n}(1)$ Elder et al. (2023a). They proved that $|\mathrm{FR}_{n}(1)|$ was equal to the number of Boolean intervals in the weak order lattice of $\mathfrak{S}_{n}$ (Elder et al., 2023a, Theorem 1.2). It remains an open problem to determine if there is some correspondence between the set $\mathrm{FR}_{n}(\ell)$ and other structures in the weak order lattice of $\mathfrak{S}_{n}$.
5 Unit interval rational parking functions and preferential arrangements
@@ 5.1 Preferential arrangements
Gross (1962) defines preferential arrangements as the number of ways $n$ distinguishable objects can be distributed into $r$ distinguishable boxes where $r$ varies $1\leq r\leq n$. Now recall that Fubini rankings are rankings of $n$ competitors in which ties are allowed (Definition 2.3). There is a bijection between preferential arrangements and Fubini rankings, which we prove next in Theorem 5.2.
@@
For surjectivity: Let $B_{1}B_{2}\cdots B_{r}\in\mathrm{PA}_{n,r}$. We construct a tuple in $[n]^{n}$ as follows: If $i\in B_{1}$, then set $a_{i}=1$. If $i\in B_{j}$ with $1<j\leq r$, then set $a_{i}=1+\sum_{k=1}^{j-1}|B_{k}|$. Let $\alpha=(a_{1},a_{2},\ldots,a_{n})$. Claim: $\alpha\in\mathrm{FR}_{n,r}$. First note that $\alpha$ has length $n$ as $\sum_{j=1}^{r}|B_{j}|=n$. By construction, the tuple $\alpha$ has precisely $r$ distinct values i.e., $r$ ranks. We need to show that the ranks in $\alpha$ satisfy the needed condition for a Fubini ranking. Note that there are $|B_{1}|$ entries in $\alpha$ with value one. Then the next rank appearing in $\alpha$ is precisely $1+|B_{1}|$, accounting for the fact that $|B_{1}|$ competitors tied for rank one. Note that then rank $1+|B_{2}|$ appears in $\alpha$ at index $i$ whenever $i\in B_{2}$. Continuing in this fashion, we then guarantee that all $r$ ranks are utilized and that no ranks are skipped. Thus $\alpha$ is a Fubini ranking, as claimed. It suffices to show that $\varphi(\alpha)=B_{1}B_{2}\cdots B_{r}$. By definition, all the elements of $\alpha$ in $I(1)$ will be placed in $B_{1}$. Similarly, if $x$ is the next available rank in $\alpha$, all the elements with rank $x$ will be placed in $B_{2}$, repeating this process for each of the $r$ available ranks ensures that $\varphi(\alpha)=B_{1}B_{2}\cdots B_{r}$, as claimed.
@@
Proof 5.5.
Given that $\mathrm{PA}_{n}=\cup_{r=1}^{n}\mathrm{PA}_{n,r}$ and $\mathrm{FR}_{n}=\cup_{r=1}^{n}\mathrm{FR}_{n,r}$, and the fact that the function $\varphi$ in Theorem, 5.2 preserves the number of distinct ranks/boxes, we have that $\varphi:\mathrm{FR}_{n}\to\mathrm{PA}_{n}$ defined analogously as in Theorem 5.2 is also a bijection.
@@
5.2 Unit interval rational parking functions and barred preferential arrangements
From Bradt et al. (2024), we know that the set of Fubini rankings are in bijection with the unit interval parking functions. In Theorem 5.2, we establish that Fubini rankings are also in bijection with preferential arrangements. Thus, preferential arrangements are in bijection with unit interval parking functions. In fact more is true. In this section, we establish that unit rational parking functions are in bijection with barred preferential arrangements.
Barred preferential arrangements were first introduced by Ahlbach et al. (2013) and have been also studied by Nkonkobe et al. (2020) and they are a generalization of preferential arrangements, which we define next.
@@ 5.2 Unit interval rational parking functions and barred preferential arrangements
Definition 5.12 (Definition 2.7, Bradt et al. (2024)).
Given a parking function $\alpha=(a_{1},\ldots,a_{n})\in\mathrm{PF}_{n}$, let $\alpha^{\prime}=(a^{\prime}_{1},\ldots,a^{\prime}_{n})$ be the weakly increasing rearrangement of $\alpha$. The partition of $\alpha^{\prime}$ as a concatenation $\alpha^{\prime}=\pi_{1}|\pi_{2}|\cdots|\pi_{m}$, where $\pi_{j}$ begins at (and includes) the $j$th entry $\alpha^{\prime}_{i}$ satisfying $\alpha^{\prime}_{i}=i$, is called the block structure of $\alpha$. Each $\pi_{i}$ is called a block of $\alpha$.
@@
Corollary 5.13 (Corollary 3.1, Elder et al. (2023a)).
Let $\alpha\in\mathrm{UPF}_{n}$ and $b_{1}\,|\,b_{2}\,|\,\cdots\,|\,b_{m}$ be its block structure. For each $j\in[m]$, let $i_{j}$ be the minimal element of $b_{j}$. Consider any $j\in[m-1]$. If $|b_{j}|=1$, then $b_{j}=(i_{j})$ and $i_{j+1}=i_{j}+1$. Otherwise, if $|b_{j}|=2$, then $b_{j}=(i_{j},i_{j})$ and $i_{j+1}=i_{j}+2$. Otherwise, $|b_{j}|\geq 3$, $b_{j}=(i_{j},i_{j},\underbrace{i_{j}+1,i_{j}+2,\ldots,i_{j}+|b_{j}|-2}_{|b_{j}|-2\mbox{ terms }})$ and $i_{j+1}=i_{j}+|b_{j}|$.
@@
$a_{i}=\begin{cases}m_{j}\mbox{if $i$ is the first entry in $B_{j}$}\\ m_{j}+y-2\mbox{if $i$ is the $y$th entry in $B_{j}$}.\end{cases}$
@@
Now we put all the steps above together. Since $\mathcal{O}(\alpha)=x(q)$, steps 1 and 2 ensure we have a barred preferential arrangement in $\mathrm{BPA}(n,m-n)$. Then we add parenthesis satisfying $|\pi_{i}|=B_{i}$ for all $i\in[r]$ and so step 3 yields precisely the parenthesization of $q$. Therefore, $\vartheta(\alpha)=q$.
We conclude with some enumerative consequences of Theorem 5.19 and the work of Ahlbach et al. (2013).
@@
(Ahlbach et al., 2013, Theorem 2) If $m\geq 1$, then
$|\mathrm{\mathrm{IPF}}_{n,m}(1)|=\frac{1}{2^{m-n}(m-n)!}\sum_{i=0}^{m-n}s(m-n+1,i+1)\thinspace|\mathrm{\mathrm{IPF}}_{n+i}(1)|$ (10)
@@
(Ahlbach et al., 2013, Theorem 3) If $m\geq 0$ and $n\geq 1$, then
$|\mathrm{\mathrm{IPF}}_{n,m}(1)|=\sum_{k=0}^{n}k!\thinspace S(n,k)\thinspace\binom{m-n+k}{k}$ (11)
@@
Note that setting $m=n$ in Theorem 11 recovers Theorem 2.2. We conclude with the following illustrative example.
@@
Acknowledgements.
This material is based upon work supported by the National Science Foundation under Grant No. DMS-1659138 and the Sloan Grant under Grant No. G-2020-12592.